How do you calculate the projected dividends when we only know the current price of the stock?
Answer #1

As per Dividend Discount Model,

P0= D1/(r-g)

Where P0= Current price of the stock, D1= Dividend expected in year 1, r= rate of return expected and g= constant growth rate of dividend.

Re-arranging the variables, Dividend next year (D1)= P0*(r-g)

Example:

Current Share price= $50, Expected rate of return= 8% and constant growth rate of dividend= 3%.

Then, Dividend next year D1 = $50*(0.08-0.03) = $50*0.05 = $2.5

Dividend for subsequent years will be higher by 3% of the previous year amount.